OpenStudy (anonymous):

After a 60% reduction, you purchase a new sofa on sale for $320. What was the original price of the sofa?

jimthompson5910 (jim_thompson5910):

let x = original price

jimthompson5910 (jim_thompson5910):

if you take 60% of some unknown value x, you get 0.60x

jimthompson5910 (jim_thompson5910):

you then subtract it from the original price x to get x - 0.60x 1x - 0.60x (1-0.60)x 0.40x

jimthompson5910 (jim_thompson5910):

so the sale price is 0.40x, but we also know the sale price is $320, so 0.40x = 320 x = ???

OpenStudy (anonymous):

$800! thank you!

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

check: 60% of 800 = 0.6*800 = 480 so the discount is $480 (ie you save $480) 800-480 = 320 so it works out